# Mantis: Interleaved Multi-Image Instruction Tuning

### v2

@@ Abstract
Large multimodal models (LMMs) have shown great results in single-image vision language tasks. However, their abilities to solve multi-image visual language tasks is yet to be improved. The existing LMMs like OpenFlamingo, Emu2, Idefics gain their multi-image ability through pre-training on hundreds of millions of noisy interleaved image-text data from the web, which is neither efficient nor effective. In this paper, we aim to build strong multi-image LMMs via instruction tuning with academic-level resources. Therefore, we meticulously construct Mantis-Instruct containing 721K multi-image instruction data to train a family of models Mantis. The instruction tuning empowers Mantis with different multi-image skills like co-reference, comparison, reasoning, and temporal understanding. We evaluate Mantis on five multi-image benchmarks and seven single-image benchmarks. Mantis-SigLIP can achieve SoTA results on all the multi-image benchmarks and beat the strongest multi-image baseline, Idefics2-8B by an average of 11 absolute points. Notably, Idefics2-8B was pre-trained on 140M interleaved multi-image data, which is 200x larger than Mantis-Instruct. We observe that Mantis performs equivalently well on the held-in and held-out benchmarks, which shows its generalization ability. Notably, we found that Mantis can even match the performance of GPT-4V on multi-image benchmarks. We further evaluate Mantis on single-image benchmarks and demonstrate that Mantis also maintains a strong single-image performance on par with CogVLM and Emu2. Our results show that multi-image abilities are not necessarily gained through massive pre-training, instead, it can be gained by the low-cost instruction tuning. Our work provides new perspectives on how to improve LMMs’ multi-image abilities.
@@ 1 Introduction
Large Multimodal Models (LMMs) have advanced significantly in recent years. Both closed-source models like GPT-4V [1], Gemini [2], Reka [3], MM1 [4] and open-source models like LLaVA [5, 6], LLaVA-NeXT [7], BLIP [8], CogVLM [9], Idefics [10] have shown strong visual-language understanding and generation capabilities in single-image tasks like VQA [11], TextVQA [12], GQA [13], etc. Despite the significant interest in improving LMMs’ performance on single-image vision-language tasks, relatively less work attempts to improve LMMs’ capability to solve multi-image vision-language tasks. To the best of our knowledge, only a few open-source models (e.g. Idefics [10], OpenFlamingo [14], Emu2 [15]) and VILA [16] support multi-image inputs.
@@
We argue that multi-image visual ability is also a crucial in real-world applications. Specifically, we categorize the ability into four skills including (1) Co-reference: understanding the references like “second image” in the natural language expression and grounding it on the referred image to generate a response. (2) Comparison: capturing the nuances and commonalities between several images. (3) Reasoning: capturing the information across multiple images and reasoning over these multiple pieces to derive the response. (4) Temporal understanding: observing multiple frames of a video or a scene to understand the temporal information like actions, behaviors, interactions, etc. These are the four major skills we want the LMMs to encompass.
@@
The existing multi-image LMMs like Idefics [10], OpenFlamingo [14], Emu2 [15], MM1 [4], VILA [16] heavily rely on pre-training on a massive interleaved image-text web documents like MMC4 [17], Obelics [10]. For example, MMC4 contains 500 million examples, while Obelics contains 140 million examples. Such a pre-training procedure would require a massive amount of computation, which hampered its adoption in most open-source LMM training pipelines. In this paper, we attempt to build LMMs with academic-level resources to support interleaved multi-image inputs. Specifically, we made a few efforts along this line:
@@
Dataset: We create the first multi-image instruction-tuning dataset Mantis-Instruct. It has a total of 721K instances, consisting of 14 subsets to cover all the multi-image skills. Among the 14 subsets, 10 subsets are from the existing datasets. For example, NLVR2 [18], IconQA [19], etc are used to cover ‘reasoning’ skill; DreamSim [20], Birds-to-Words [21], etc are used to cover ‘comparison’ skill; NExT-QA [22], STAR [23], etc are used to cover ‘temporal understanding’ skill. We additionally curate four new datasets LLaVA-665k-multi, LRV-multi to cover the ‘coref’ skill and Contrast-Caption, Multi-VQA to broaden the ‘reasoning’ skill.
@@
Architecture: Similar to LLaVA [7], we start from a pre-trained language model like LLaMA-3 [24] or Fuyu [25] and a vision transformer encoder from CLIP [26] or SigLIP [27]. We use the same multimodal projector as LLaVA to map the vision embeddings to the text embeddings space to simply concatenate them with the text embeddings. We also define a text-image interleaving format: “…(image {i}: <BOI>image embeddings</EOI>)…”, where <BOI> and</EOI> are the image delimiters. We explored multiple instantiations and found this optimal design.
@@
We train Mantis by combining Mantis-Instruct with another 268K single-image vision-language data to balance the multi-image and single-image abilities. To evaluate the multi-image skills, we adopt four existing benchmarks NLVR2 [18], QBench [28], BLINK [29], MVBench [30] and Mantis-Eval to cover all the mentioned skills. We include 15 competitive baseline models including single-image LMMs like LLaVA-Next [7], Qwen-VL [31] and multi-image LMMs like GPT-4V, Idefics1/2, etc. For single-image LMMs, we merge multiple images horizontally as a single-image input. By training on 16 x A100-40G for 36 hours, Mantis achieves state-of-the-art performance on all five multi-image tasks, excelling on all the multi-image skills. Notably, Mantis outperforms the strongest baseline Idefics2 [10] by an average of 11 absolute points. The held-in and held-out results are equivalently strong, which shows the generalization ability of Mantis. These results are encouraging because we show that the low-cost instruction tuning on 721K high-quality data can lead to a much better generalization performance than other LMMs intensively pre-trained on 100x larger datasets. Mantis also preserves its strong performance of single-image tasks and is on par with CogVLM [9] and Emu2 [15]. We also present insights on the architecture selection and inference approaches based on the results in  section 3. We believe Mantis-Instruct can serve as an essential baseline for future studies on multi-image LMMs due to its simplicity and low cost. We will release all the code, data, and models to help with the reproducibility of our results.
@@ 2.1 Method
Most existing LMMs do not support multi-image inputs, either due to the architecture design [32, 8], or the lack of associated code support [33, 25]. To enable multi-image training and inference, we have modified the LLaVA architecture and added code support for it. Idefics2 [34] already supports multi-image as inputs naturally, we thus inherit their structure.
@@
Text-Image Interleaving Format:
A proper text-image interleaving format can make an LLM easier to acquire multi-image understanding and reasoning ability. We contend that a good text-image interleaving format should: (1) mark boundaries between images clearly, and (2) denote the serial number of images. Following this principle, we designed our interleaving format as follows: "(image {i}: <BOI><image><EOI>)", where <BOI> is the begin of image token and <EOI> is the end of image token. <image> is the placeholder for image patches. This format adds clear separators between images, and gives LMM serial information of the image through "image {i}". In practice, we set <BOI> and <EOI> to be <Image> and </Image> respectively. Previous work also demonstrates the effectiveness of this approach through a comprehensive ablation study [35].
@@ 2.2 Mantis-Instruct: A large scale multiple image question answering dataset
Table 1: The statistics of Mantis-Instruct. We list the number of examples, average images per example, maximum image per example, average turns, average text token lengths, and average text+image token lengths. $\clubsuit$ means the new subsets we constructed in this paper. Here we use Llava’s tokenizer and assume each image occupies 576 image tokens, which is the number of tokens for each image. Here "Avg LengthT" denotes the average number of tokens for the text part, while "Avg LengthT&I" means the average number of tokens of text and images in total.
@@
requires the model to further make inferences by combining its world knowledge with the information collected using the co-reference and comparison ability. It distinguishes from the comparison skill by introducing more complex human instruction associated with the images. We gather NLVR2, IconQA, Contrast-Caption by reformatting captioning datasets, ImageCoDe, and our self-collected Multi-VQA to further equip the model with reasoning ability. The topics include logical reasoning, counting, image matching, image retrieval, and free-form multi-image QA.
@@
Figure 2: Illustrations of Mantis-Instruct datasets from Contrast-Captioning and Multi-VQA.
@@ 3 Experiments
In this section, we aim to evaluate Mantis’s ability on both multi-image and single-image vision-language tasks. For multi-image tasks, depending on whether the LMM can handle multiple images, we either concatenate multiple images (‘merge’) horizontally as a single image or feed multiple images as a ‘sequence’ to the model. We have trained four model variants listed in  Table 2, where we list the architecture and dataset details. Mantis-CLIP and Mantis-SigLIP only adopt the CC3M 560K subset for feature alignment. Mantis-Flamingo and Mantis-Idefics2 are initialized from OpenFlamingo and Idefics2, respectively, which have been intensively pre-trained on the multi-image corpus. After this pre-training and preparation, we fine-tune them on Mantis-Instruct to get Mantis. Please refer to  subsection A.3 for details of hyper-parameters and settings.
@@
Mantis-Flamingo CLIP MPT-7B MMC4, LAION 2421M Mantis-Instruct 576
@@ 3.1 Baselines
For ‘merge’ evaluation, we include BLIP-2-13B [8], InstructBLIP-13B [36], Qwen-VL-Chat [31], LLaVA-1.5-7B [5], LLaVA-1.6-7B [7], Kosmos2 [33], and Fuyu-8B [25]. For ‘sequence’ evaluation, we include OpenFlamingo [14], Otter [37], VideoLLaVA [38], Emu2-Chat-34B [15], VILA [39], Idefics1 [10], Idefics2 [10], GPT-4V [1]. Additionally, for single-image tasks, we include InstructBLIP-7B-Vicuna [36], Yi-VL-6B [40].
@@ 3.2 Evaluation Benchmarks
For multi-image tasks, we use 2 held-in benchmarks: NLVR2 [18] and Qbench [28]; and 3 held-out benchmarks: Mantis-Eval, BLINK [29], and MVBench [30]. We report detailed statistics in  Table 3.
NLVR2 [18] evaluates the model’s ability to conduct logical reasoning across the contents of images. It asks the model to compare the contents of the two given images and judge whether a given statement is correct or not. All questions are multiple-choice. We use the test-public split for evaluation.
Qbench [28] is a benchmark evaluating whether LMMs can properly judge and compare the quality of a benchmark. Q-bench aims to evaluate the low-level visual abilities of LMMs, such as judging the image quality. All questions are multiple-choice. In our experiments, we evaluate the Qbench2-A2-pair dev set, where a low-level visual question is asked based on multiple image contents.
@@
BLINK [29] is a benchmark on core visual perception abilities, where most of the tasks can be solved by humans “within a blink” (e.g., relative depth estimation, visual correspondence, forensics detection, and multi-view reasoning). Some questions in the benchmark involved multiple images, such as image similarity comparison. We report results in the validation set of the benchmark.
MVBench [30] is a comprehensive multimodal video understanding benchmark covering 20 challenging video tasks. The questions cannot be solved with a single frame, which necessitates understanding an image sequence to solve the questions in the benchmark. We report results in test split.
Table 3: Statistics of the evaluation dataset. NLVR2 and Mantis-Eval are held-in dataset, while the rest are held-out. For MVBench, we extract 8 frames per video for the evaluation.
@@
Table 4: Evaluation results on multi-image benchmarks. The highest score of open-source models for each tasks is bolded. The highest score of the open-source baseline model is underscored. $\Delta$ is the performance gains of Mantis compared to the best baseline model for each task.
Models Size Format NLVR2 Q-Bench Mantis-Eval BLINK MVBench Avg
GPT-4V - sequence 88.80 76.52 62.67 51.14 43.50 64.5
Open Source Models
Random - - 48.93 40.20 23.04 38.09 27.30 35.5
Kosmos2 1.6B merge 49.00 35.10 30.41 37.50 21.62 34.7
LLaVA-v1.5 7B merge 53.88 49.32 31.34 37.13 36.00 41.5
LLava-V1.6 7B merge 58.88 54.80 45.62 39.55 40.90 48.0
@@
Fuyu 8B merge 51.10 49.15 27.19 36.59 30.20 38.8
BLIP-2 13B merge 59.42 51.20 49.77 39.45 31.40 46.2
InstructBLIP 13B merge 60.26 44.30 45.62 42.24 32.50 45.0
CogVLM 17B merge 58.58 53.20 45.16 41.54 37.30 47.2
OpenFlamingo 9B sequence 36.41 19.60 12.44 39.18 7.90 23.1
Otter-Image 9B sequence 49.15 17.50 14.29 36.26 15.30 26.5
Idefics1 9B sequence 54.63 30.60 28.11 24.69 26.42 32.9
VideoLLaVA 7B sequence 56.48 45.70 35.94 38.92 44.30 44.3
Emu2-Chat 37B sequence 58.16 50.05 37.79 36.20 39.72 44.4
@@
Mantis-CLIP 8B sequence 84.66 66.00 55.76 47.06 48.30 60.4
Mantis-SigLIP 8B sequence 87.43 69.90 59.45 46.35 50.15 62.7
Mantis-Flamingo 9B sequence 52.96 46.80 32.72 38.00 40.83 42.3
Mantis-Idefics2 8B sequence 89.71 75.20 57.14 49.05 51.38 64.5
$\Delta$ over SOTA - - +2.84 +18.20 +8.30 +3.87 +1.98 +11.0

### v3

@@ Abstract
Large multimodal models (LMMs) have shown great results in single-image vision language tasks. However, their abilities to solve multi-image visual language tasks is yet to be improved. The existing LMMs like OpenFlamingo, Emu2, and Idefics gain their multi-image ability through pre-training on hundreds of millions of noisy interleaved image-text data from the web, which is neither efficient nor effective. In this paper, we aim to build strong multi-image LMMs via instruction tuning with academic-level resources. Therefore, we meticulously construct Mantis-Instruct containing 721K multi-image instruction data to train a family of Mantis models. The instruction tuning empowers Mantis with different multi-image skills like co-reference, comparison, reasoning, and temporal understanding. We evaluate Mantis on 8 multi-image benchmarks and 6 single-image benchmarks. Mantis-Idefics2 can achieve SoTA results on all the multi-image benchmarks and beat the strongest multi-image baseline, Idefics2-8B by an average of 13 absolute points. Notably, Idefics2-8B was pre-trained on 140M interleaved multi-image data, which is 200x larger than Mantis-Instruct. We observe that Mantis performs equivalently well on the held-in and held-out benchmarks, which shows its generalization ability. We further evaluate Mantis on single-image benchmarks and demonstrate that Mantis also maintains a strong single-image performance on par with CogVLM and Emu2. Our results show that multi-image abilities are not necessarily gained through massive pre-training, instead, they can be gained by low-cost instruction tuning. The training and evaluation of Mantis has paved the road for future work to improve LMMs’ multi-image abilities.
@@ 1 Introduction
Large Multimodal Models (LMMs) have advanced significantly in recent years. Both closed-source models like GPT-4V (Achiam et al., 2023), Gemini (Team et al., 2023), Reka (Ormazabal et al., 2024), MM1 (McKinzie et al., 2024) and open-source models like LLaVA (Liu et al., 2023c; b), LLaVA-NeXT (Liu et al., 2024), BLIP (Li et al., 2023c), CogVLM (Wang et al., 2023), Idefics (Laurençon et al., 2023) have shown strong visual-language understanding and generation capabilities in single-image tasks like VQA (Antol et al., 2015), TextVQA (Singh et al., 2019b), GQA (Hudson & Manning, 2019), etc. Despite the significant interest in improving LMMs’ performance on single-image vision-language tasks, relatively less work attempts to improve LMMs’ capability to solve multi-image vision-language tasks. To the best of our knowledge, only a few open-source models (e.g. Idefics (Laurençon et al., 2023), OpenFlamingo (Awadalla et al., 2023), Emu2 (Sun et al., 2023)) and VILA (Lin et al., 2023c) support multi-image inputs.
@@
We argue that multi-image visual ability is also a crucial in real-world applications. Specifically, we categorize the ability into four major skills we want the LMM to encompass: (1) Co-reference: understanding the references like “second image” in the natural language expression and grounding it on the referred image to generate a response. (2) Comparison: capturing the nuances and commonalities between several images. (3) Reasoning: capturing the information across multiple images and reasoning over these multiple pieces to derive the response. (4) Temporal understanding: observing multiple frames of a video or a scene to understand the temporal information like actions, behaviors, interactions, etc.
@@
The existing multi-image LMMs like Idefics (Laurençon et al., 2023), OpenFlamingo (Awadalla et al., 2023), Emu2 (Sun et al., 2023), MM1 (McKinzie et al., 2024), VILA (Lin et al., 2023c) heavily rely on pre-training on a massive interleaved image-text web documents like MMC4 (Zhu et al., 2024), Obelics (Laurençon et al., 2023). For example, MMC4 contains 500 million examples, while Obelics contains 140 million examples. Such a pre-training procedure would require a massive amount of computation, which hampered its adoption in most open-source LMM training pipelines. In this paper, we attempt to build LMMs with academic-level resources to support interleaved multi-image inputs. Specifically, we made a few efforts along this line:
@@
Dataset: We create the first multi-image instruction-tuning dataset Mantis-Instruct. It has a total of 721K instances, consisting of 14 subsets to cover all the multi-image skills. Among the 14 subsets, 10 subsets are from the existing datasets. For example, NLVR2 (Suhr et al., 2018), IconQA (Jhamtani & Berg-Kirkpatrick, 2018), etc are used to cover ‘reasoning’ skill; DreamSim (Fu et al., 2023), Birds-to-Words (Forbes et al., 2019), etc are used to cover ‘comparison’ skill; NExT-QA (Xiao et al., 2021), STAR (Wu & Yu, 2021), etc are used to cover ‘temporal understanding’ skill. Additionally, 4 new subsets are newly curated, where LLaVA-665k-multi, and LRV-multi are used to to cover the ‘coref’ skill, Contrast-Caption, and Multi-VQA are used to broaden the ‘reasoning’ skill. The interleaving image placeholders are inserted into the text under various heuristics.
@@
Architecture: Similar to LLaVA (Liu et al., 2024), we start from a pre-trained language model like LLaMA-3 (AI@Meta, 2024) or Fuyu (Adept AI, 2023) and a vision transformer encoder from CLIP (Radford et al., 2021) or SigLIP (Zhai et al., 2023). We use the same multimodal projector as LLaVA to map the vision embeddings to the text embeddings space to simply concatenate them with the text embeddings. We also define a text-image interleaving format: “…(image {i}: <BOI>image embeddings</EOI>)…”, where <BOI> and</EOI> are the image delimiters. We explored multiple instantiations and found this optimal design.
@@
We train Mantis by combining Mantis-Instruct with another 268K single-image vision-language data to balance the multi-image and single-image abilities. To evaluate the multi-image skills, we adopt 8 benchmarks including NLVR2 (Suhr et al., 2018), BLINK (Fu et al., 2024), MVBench (Li et al., 2023d), Mantis-Eval, etc. to cover all the mentioned skills. We include 15 competitive baseline models including single-image LMMs like LLaVA-Next (Liu et al., 2024), Qwen-VL (Bai et al., 2023) and multi-image LMMs like GPT-4V, Idefics1/2, etc. For single-image LMMs, we merge multiple images horizontally as a single-image input. By training on 16 x A100 for 36 hours, Mantis achieves state-of-the-art performance on all the multi-image tasks, while preserving competitive performance in single-image tasks. Notably, Mantis-Idefics2 outperforms the strongest baseline Idefics2 (Laurençon et al., 2023) by an average of 13 absolute points on the 7 avaialble benchmarks (excluding MileBench), where Idefics2 is multi-image pre-trained, showcasing the effectiveness of multi-image instruction-tuning. The held-in and held-out results are equivalently strong, which shows the generalization ability of Mantis.
These results are encouraging because we show that the low-cost instruction tuning on 721K high-quality data can lead to a much better generalization performance than other LMMs intensively pre-trained on 100x larger datasets. Mantis also preserves its strong performance of single-image tasks and is on par with CogVLM (Wang et al., 2023) and Emu2 (Sun et al., 2023). We also present insights on the architecture selection and inference approaches based on the results in  section 3. We believe Mantis-Instruct can serve as an essential baseline for future studies on multi-image LMMs due to its simplicity and low cost. We will release all the code, data, and models to help with the reproducibility of our results.
@@ 2.1 Method
Most existing LMMs do not support multi-image inputs, either due to the architecture design (Huang et al., 2016; Li et al., 2023c), or the lack of associated code support (Peng et al., 2023; Adept AI, 2023). To enable multi-image training and inference, we have modified the LLaVA architecture and added code support for it. Idefics2 (Laurenccon et al., 2024) already supports multi-image as inputs naturally, we thus inherit their structure.
@@
Interleaving Text-Image:
A proper text-image interleaving format can help acquire multi-image understanding and reasoning ability. We contend that a good text-image interleaving format should: (1) mark boundaries between images clearly, and (2) denote the serial number of images. Following this principle, we designed our interleaving format as follows: "(image {i}: <BOI><image><EOI>)", where <BOI> is the begin of image token and <EOI> is the end of image token. <image> is the placeholder for image patches. This format adds clear separators between images, and gives serialized information of the image through "image {i}". In practice, we set <BOI> and <EOI> to be <Image> and </Image> respectively. Previous work also demonstrates its effectiveness through a comprehensive ablation study (Wu et al., 2024).
@@ 2.2 Mantis-Instruct: A large scale multiple image question answering dataset
Table 1: The statistics of Mantis-Instruct. We list the number of examples, average images per example, maximum image per example, average turns, average text token lengths, and average text+image token lengths. $\clubsuit$ means the new subsets we constructed in this paper. Here we use Llava’s tokenizer and assume each image occupies 576 image tokens, which is the number of tokens for each image. Here "Avg LengthT" denotes the average number of tokens for the text part, while "Avg LengthT&I" means the average number of tokens of text and images in total. Heuristics about how each subset is curated can be found in  subsection A.1.
Output format:
Question: First question?
Answer: The answer to the first question.
Question: Second question?
Answer: The answer to the second question.
Table 2: Propmt template used to curate the Multi-VQA subset.
@@
requires the model to further make inferences by combining its world knowledge with the information collected using the co-reference and comparison ability. It distinguishes from the comparison skill by introducing more complex human instruction associated with the images. We gather NLVR2, IconQA, Contrast-Caption by reformatting captioning datasets, ImageCoDe, and our self-collected Multi-VQA to further equip the model with reasoning ability. The topics include logical reasoning, counting, image matching, image retrieval, and free-form multi-image QA. Multi-VQA is synthesized by prompting GPT-4V, where the prompt template is shown in  Table 2.
@@
2.3 Heuristics of data curation
Similar to visual instruction tuning (Liu et al., 2023c), data curation is an important step to ensure good performance. During the curation, we found some heuristics that could improve performance. We list them below. Note that the first heuristics is a simple way to diversify QA format to avoid overfitting, and the rest two heuristics have already been proven effective by previous work Co-Instruct (Wu et al., 2024). We thus adopt them directly to avoid duplicate efforts to confirm their effectiveness.
  1. 1.
    Conflicts between short answers and long answers. Many datasets we collected are domain-specific, and some of them are classification tasks like NLVR2 (Suhr et al., 2018) and DreamSim (Fu et al., 2023). For these datasets, we convert them into multiple-choice QA format to avoid conflicts between repeated short-answers and some long answers from other datasets.
  2. 2.
  3. 3.
    Positions of the "<image>" placeholder. For each item with multiple images, we write some rules to randomly put the image placeholders at the beginning of the first question or the end of the first question. This technique is also used by the curation LLaVA-665k (Liu et al., 2023b).
Figure 2: Illustrations of Mantis-Instruct datasets from Contrast-Captioning and Multi-VQA. Contrastive-Captioning is constructed from existing image captioning datasets to enhance LMM’s multi-image ‘co-reference’ ability. Multi-VQA is GPT-4V synthesized multi-turn multi-image data used to enhance LMM’s multi-image ‘reason’ ability.
@@ 3 Experiments
In this section, we aim to evaluate Mantis’s ability on both multi-image and single-image vision-language tasks. For multi-image tasks, depending on whether the LMM can handle multiple images, we either concatenate multiple images (‘merge’) horizontally as a single image or feed multiple images as a ‘sequence’ to the model. We have trained four model variants listed in  Table 3, where we list the architecture and dataset details. Mantis-CLIP and Mantis-SigLIP only adopt the CC3M 560K subset for feature alignment. Mantis-Flamingo and Mantis-Idefics2 are initialized from OpenFlamingo and Idefics2, respectively, which have been intensively pre-trained on the multi-image corpus. We also include Mantis-Fuyu to investigate the significance of vision encoder in the LMM architecture. After this pre-training and preparation, we fine-tune them on Mantis-Instruct to get Mantis.
@@
Mantis-Fuyu - Persimmon-8B Unknown Unknown Mantis-Instruct [1,2304]
Mantis-Flamingo CLIP MPT-7B MMC4, LAION 2421M Mantis-Instruct 576
@@
3.1 Training Details
For Mantis-CLIP and Mantis-SigLIP, we first pre-train the multimodal projector on LLaVA pre-train data, where the learning rate is set to 1e-3, the batch size is set to 256. After pre-training, we fine-tune the 2 models on the Mantis-Instruct. For Mantis-Fuyu, Mantis-Flamingo and Mantis-Idefics2, we directly fine-tune them on Mantis-Instruct since they have already been pre-trained on millions of data.
During the fine-tuning, we train each model on the data for 1 epoch, with a batch size of 128. The maximum context length is set to 8192. The learning rate is all set to 1e-5, except for Idefics2, where the learning rate is set to 5e-6 to better preserve its original knowledge. We set the warmup ratio to be 0.03 and use a cosine learning rate scheduler. We speed up our training and inference with Flash-attention2 (Dao, 2023). We use DeepSpeed Zero-3 (Aminabadi et al., 2022) for full fine-tuning. We apply QLoRA (Dettmers et al., 2023) along with DoRA (yang Liu et al., 2024) to more efficiently do comprehensive ablation studies under limited resources. All full fine-tuning ran on 16 A100 GPUs while the ablation study ran on 8 A100 GPUs.
@@ 3.2 Baselines
For ‘merge’ evaluation, we include BLIP-2-13B (Li et al., 2023c), InstructBLIP-13B (Dai et al., 2023), Qwen-VL-Chat (Bai et al., 2023), LLaVA-1.5-7B (Liu et al., 2023c), LLaVA-1.6-7B (Liu et al., 2024), Kosmos2 (Peng et al., 2023), and Fuyu-8B (Adept AI, 2023). For ‘sequence’ evaluation, we include OpenFlamingo (Awadalla et al., 2023), Otter (Li et al., 2023b), VideoLLaVA (Lin et al., 2023a), Emu2-Chat-34B (Sun et al., 2023), VILA (Lin et al., 2023b), Idefics1 (Laurençon et al., 2023), Idefics2 (Laurençon et al., 2023), GPT-4V (Achiam et al., 2023). Additionally, for single-image tasks, we include InstructBLIP-7B-Vicuna (Dai et al., 2023), Yi-VL-6B (AI et al., 2024).
@@ 3.3 Evaluation Benchmarks
For multi-image tasks, we use 2 held-in benchmarks: NLVR2 (Suhr et al., 2018) and Qbench (Wu et al., 2023a); and 3 held-out benchmarks: Mantis-Eval, BLINK (Fu et al., 2024), and MVBench (Li et al., 2023d). We report detailed statistics in  Table 4.
NLVR2 (Suhr et al., 2018) evaluates the model’s ability to conduct logical reasoning across the contents of images. It asks the model to compare the contents of the two given images and judge whether a given statement is correct or not. All questions are multiple-choice. We use the test-public split for evaluation.
Qbench (Wu et al., 2023a) is a benchmark evaluating whether LMMs can properly judge and compare the quality of a benchmark. Q-bench aims to evaluate the low-level visual abilities of LMMs, such as judging the image quality. All questions are multiple-choice. In our experiments, we evaluate the Qbench2-A2-pair dev set, where a low-level visual question is asked based on multiple image contents.
@@
BLINK (Fu et al., 2024) is a benchmark on core visual perception abilities, where humans can solve most tasks “within a blink” (e.g., relative depth estimation, visual correspondence, forensics detection, and multi-view reasoning). Some questions in the benchmark involved multiple images, such as image similarity comparison. We report results in the validation set of the benchmark.
MVBench (Li et al., 2023d) is a comprehensive multimodal video understanding benchmark covering 20 challenging video tasks. The questions cannot be solved with a single frame, necessitating understanding an image sequence to solve the questions in the benchmark. We extract 8 frames per video for the evaluation. We report results in test split.
MileBench (Song et al., 2024) is a novel benchmark designed to test the multimodal long-context capabilities of MLLMs. It comprises realistic evaluation and diagnostic evaluation. The former involves tasks like temporal understanding and semantic multi-image comprehension. The later diagnostic evaluation involves TextNeedle and ImageNeedle retrieval tasks. We only reported the performance of the realistic evaluation. The results of the diagnostic evaluation can be found in MileBench’s paper.
MuirBench (Wang et al., 2024a) is a comprehensive benchmark consisting of 12 diverse multi-image tasks, such as scene understanding, ordering, etc. It contains 2,600 multiple-choice questions with 11,264 images involved in total. We report the overall average performance across the 12 tasks.
MMIU (Meng et al., 2024) is a multi-image evaluation benchmark encompassing 7 types of multi-image relationships, 52 tasks, 77K images, and 11K multiple-choice questions. We report the overall average performance across all the tasks.
@@
MileBench Mixed MCQ 6,440 15.2 109 9,294
MuirBench Mixed MCQ 2,600 4.3 9 2,502
MMIU Mixed MCQ 11,600 6.6 32 3,852
Table 4: Statistics of the evaluation dataset. Held-in benchmarks, NLVR2 and Q-Bench, are marked with $\dagger$ notation. The rest of them are held-out benchmarks. “Mixed” means a .
Models Size NLVR2 Q-Bench Mantis-Eval BLINK MVBench MileBench MuirBench MMIU
GPT-4V/o - 88.80 76.52 62.67 51.14 43.50 53.00 68.00 55.70
Open Source Models (merge)
Random - 48.93 40.20 23.04 38.09 27.30 22.30 23.99 27.40
Kosmos2 1.6B 49.00 35.10 30.41 37.50 21.62 - - -
LLaVA-v1.5 7B 53.88 49.32 31.34 37.13 36.00 38.00 23.46 19.20
LLaVA-V1.6 7B 58.88 54.80 45.62 39.55 40.90 38.10 - 22.20
@@
Fuyu 8B 51.10 49.15 27.19 36.59 30.20 - - -
BLIP-2 13B 59.42 51.20 49.77 39.45 31.40 - - -
InstructBLIP 13B 60.26 44.30 45.62 42.24 32.50 - - -
CogVLM 17B 58.58 53.20 45.16 41.54 37.30 - 20.85 -
Open Source Models (sequence)
OpenFlamingo 9B 36.41 19.60 12.44 39.18 7.90 27.40 23.73 -
Otter-Image 9B 49.15 17.50 14.29 36.26 15.30 - - -
[…]
VideoLLaVA 7B 56.48 45.70 35.94 38.92 44.30 - - -
Emu2-Chat 37B 58.16 50.05 37.79 36.20 39.72 - - -
@@
Mantis-Fuyu 8B 72.41 54.10 37.79 39.87 39.12 - - -
Mantis-Flamingo 9B 52.96 46.80 32.72 38.00 40.83 - - -
Mantis-CLIP 8B 84.66 66.00 55.76 47.06 48.30 - 37.38 -
Mantis-SIGLIP 8B 87.43 69.90 59.45 46.35 50.15 47.5 36.12 -
Mantis-Idefics2 8B 89.71 75.20 57.14 49.05 51.38 - 44.50 45.60
$\Delta$ over SOTA - 2.84 18.20 8.30 3.87 1.98 3.10 9.07 17.80
Table 5: Evaluation results on multi-image benchmarks. The highest score of open-source models for each task is bolded. The highest score of the open-source baseline model is underscored. $\Delta$ is the performance gains of Mantis compared to the best baseline model for each task. We report the GPT-4o’s results on MuirBench and MMIU, and GPT-4V’s on the rest of them.